# A stochastic network approach to clustering and visualising single-cell genomic count data

### v3

@@ Abstract
Important tasks in the study of genomic data include the identification of groups of similar cells (for example by clustering), and visualisation of data summaries (for example by dimensional reduction). In this paper, we propose a novel approach to studying single-cell genomic data, by modelling the observed genomic data count matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ as a bipartite network with multi-edges. Utilising this first-principles network representation of the raw data, we propose clustering single cells in a suitably identified d-dimensional Laplacian Eigenspace (LE) via a Gaussian mixture model (GMM-LE), and employing UMAP to non-linearly project the LE to two dimensions for visualisation (UMAP-LE). This LE representation of the data estimates transformed latent positions (of genes and cells), under a latent position model of nodes in a bipartite stochastic network. We demonstrate how these estimated latent positions can enable fine-grained clustering and visualisation of single-cell genomic data, by application to data from three recent genomics studies in different biological contexts. In each data application, clusters of cells independently learned by our proposed methodology are found to correspond to cells expressing specific marker genes that were independently defined by domain experts. In this validation setting, our proposed clustering methodology outperforms the industry-standard for these data. Furthermore, we validate components of the LE decomposition of the data by contrasting healthy cells from normal and at-risk groups in a machine-learning model, thereby generating an LE cancer biomarker that significantly predicts long-term patient survival outcome in an independent validation dataset.
@@ Introduction
The structure of this papers is as follows. In Section 2, we present our novel first-principles representation of single-cell genomic count data, and our improved method of clustering single cells that we call ‘GMM-LE’ clustering, and related visualisations that we call ‘UMAP-LE’. Then in Section 3, we present the results of applying GMM-LE clustering and UMAP-LE visualisation to example data-sets relevant to human cortical development, to human embryonic development, and to a population at risk of breast cancer, and we develop and validate a machine-learning cancer biomarker from the LE representation of the data.
@@ Methods
Our approach contrasts with other recent approaches such as [modell2022spectral], in which nodes of both types are included in one unified clustering procedure, by spectral embedding of a symmetric adjacency matrix that has all nodes of all types on both margins as $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}^{\top}\\ \mathbf{A}&\mathbf{0}\end{array}\right],$ where $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ is our asymmetric adjacency matrix $\mathbf{A}=\mathbf{X}$, and $\mathbf{0}$ is an appropriately-sized matrix of 0s. This symmetric embedding leads to pairs of reflected eigenvectors (Lemma 5 of [modell2022spectral]) and allows negative eigenvalues to be found which indicate disassortative features [khor2010concurrency], which in our context allow separation of the two different types of nodes (genes and cells). For example, consider the 2-node bipartite network represented (in our setting) by the adjacency matrix $A=[1]$, i.e., the corresponding data matrix $X=[1]$ represents 1 count of 1 gene in 1 cell. Then, the symmetric embedding of $A$ described above would be $\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$ which has eigenvalues 1 and -1, with corresponding eigenvectors $[1/\sqrt{2},1/\sqrt{2}]^{\top}$ and $[-1/\sqrt{2},1/\sqrt{2}]^{\top}$. The second of these eigenvectors is disassortative, because dividing the nodes based on a separating point anywhere in $\left(-1/\sqrt{2},1/\sqrt{2}\right)$ will separate the nodes into two groups according to their types ‘cell’ and ‘gene’ as a ‘cut’ based on this second eigenvector [shi2000normalized, von2007tutorial]. Whereas there is no separating point for which the first eigenvector can be used to separate the nodes into groups.
@@ Results
Figure 7: LE cancer biomarker inference, for the breast cancer at-risk data-set. (a) A generalised linear model (with logit link) was fitted to the top 25 RSVs (right singular vectors) as covariates (predictors), with BRCA1 mutation status of the donor of the cell sample as the response; the 12th RSV of the LE is most strongly associated with BRCA1 mutation status, as a surrogate measure of predisposition to aggressive cancer. (b) A Kaplan-Meier curve shows significant association of long-term patient survival outcome with a polygenic score (PGS) derived from the 12th LSV (left singular vector) of the LE.
@@ Conclusions
The breast cancer survival outcome dataset is available from the European Genome-Phenome Archive (EGA) under accession number EGAS00000000083, from URL:
@@
Gene expression data for 24368 genes and 1904 patient volunteers were downloaded, together with patient survival outcome. There were 10 missing values in the data-set, which were replaced with K-NN imputation, with $K=5$.
@@ Acknowledgements
The work carried out by TE Bartlett was partly supported by MRC grant (MR/P014070/1). The authors acknowledge the use of the UCL High Performance Computing Facilities, and associated support services, in the completion of this work.

### v4

@@ Abstract
Important tasks in the study of genomic data include the identification of groups of similar cells (for example by clustering), and visualisation of data summaries (for example by dimensional reduction). In this paper, we propose a novel approach to studying single-cell genomic data, by modelling the observed genomic data count matrix $\mathbf{X}\in\mathbb{Z}_{\geq 0}^{p\times n}$ as a bipartite network with multi-edges. Utilising this first-principles network representation of the raw data, we propose clustering single cells in a suitably identified d-dimensional Laplacian Eigenspace (LE) via a Gaussian mixture model (GMM-LE), and employing UMAP to non-linearly project the LE to two dimensions for visualisation (UMAP-LE). This LE representation of the data estimates transformed latent positions (of genes and cells), under a latent position model of nodes in a bipartite stochastic network. We demonstrate how these estimated latent positions can enable fine-grained clustering and visualisation of single-cell genomic data, by application to data from three recent genomics studies in different biological contexts. In each data application, clusters of cells independently learned by our proposed methodology are found to correspond to cells expressing specific marker genes that were independently defined by domain experts. In this validation setting, our proposed clustering methodology outperforms the industry-standard for these data. Furthermore, we validate components of the LE decomposition of the data by contrasting healthy cells from normal and at-risk groups in a machine-learning model, thereby identifying an LE cancer biomarker that significantly predicts long-term patient survival outcome in two independent validation cohorts with data from 1904 and 1091 individuals.
@@ Introduction
The structure of this papers is as follows. In Section 2, we present our novel first-principles representation of single-cell genomic count data, and our improved method of clustering single cells that we call ‘GMM-LE’ clustering, and related visualisations that we call ‘UMAP-LE’. Then in Section 3, we present the results of applying GMM-LE clustering and UMAP-LE visualisation to example data-sets relevant to human cortical development, to human embryonic development, and to a population at risk of breast cancer, and we identify and validate a machine-learning cancer biomarker from the LE representation of the data.
@@ Methods
Our approach contrasts with other recent approaches such as [modell2022spectral], in which nodes of both types are included in one unified clustering procedure, by spectral embedding of a symmetric adjacency matrix that has all nodes of all types on both margins as $\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{A}^{\top}\\ \mathbf{A}&\mathbf{0}\end{array}\right],$ where $\mathbf{A}\in\mathbb{Z}_{\geq 0}^{p\times n}$ is our asymmetric adjacency matrix $\mathbf{A}=\mathbf{X}$, and $\mathbf{0}$ is an appropriately-sized matrix of 0s. Such a matrix always has eigenvalues in pairs of equal magnitude and opposite sign. Furthermore this symmetric embedding leads to corresponding pairs of reflected eigenvectors (Lemma 5 of [modell2022spectral]), and negative eigenvalues are known to indicate disassortative features [khor2010concurrency], which in our context allow separation of the two different types of nodes (genes and cells). For example, consider the 2-node bipartite network represented (in our setting) by the adjacency matrix $A=[1]$, i.e., the corresponding data matrix $X=[1]$ represents 1 count of 1 gene in 1 cell. Then, the symmetric embedding of $A$ described above would be $\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$ which has eigenvalues 1 and -1, with corresponding eigenvectors $[1/\sqrt{2},1/\sqrt{2}]^{\top}$ and $[-1/\sqrt{2},1/\sqrt{2}]^{\top}$. The second of these eigenvectors is disassortative, because dividing the nodes based on a separating point anywhere in $\left(-1/\sqrt{2},1/\sqrt{2}\right)$ will separate the nodes into two groups according to their types ‘cell’ and ‘gene’ as a ‘cut’ based on this second eigenvector [shi2000normalized, von2007tutorial]. Whereas there is no separating point for which the first eigenvector can be used to separate the nodes into groups.
@@ Results
Figure 7: LE cancer biomarker inference, for the breast cancer at-risk data-set. (a) A generalised linear model (with logit link) was fitted to the top 25 RSVs (right singular vectors) as covariates (predictors), with BRCA1 mutation status of the donor of the cell sample as the response; the 12th RSV of the LE is most strongly associated with BRCA1 mutation status, as a surrogate measure of predisposition to aggressive cancer. (b) and (c) Kaplan-Meier curves show significant prediction of long-term patient survival outcome by a polygenic score (PGS) derived from the 12th LSV (left singular vector) of the LE, respectively in the independent METABRIC [curtis2012genomic] and TCGA-BRCA [cancer2012comprehensive] validation datasets.
@@ Conclusions
The METABRIC breast cancer validation dataset is available from the European Genome-Phenome Archive (EGA) under accession number EGAS00000000083, from URL:
@@
Gene expression data for 24368 genes and 1904 patient volunteers were downloaded, together with patient survival outcome data for all volunteers in the study. There were 10 missing values in the genomic data-set, which were replaced with K-NN imputation, with $K=5$.
The TCGA-BRCA breast cancer validation dataset is available from the NIH GDC Data Portal under project TCGA, from URL: https://portal.gdc.cancer.gov
Gene expression data for 20499 transcripts and 1091 patient volunteers were downloaded, together with patient survival outcome data for all volunteers in the study. There were no missing values in the data-set.
@@ Acknowledgements
The work carried out by TE Bartlett was partly supported by MRC grant (MR/P014070/1). The authors acknowledge the use of the UCL High Performance Computing Facilities, and associated support services, in the completion of this work. The results published herein are in whole or in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga